I definately need help with this problem...thanks so very much
Apply the quadratic formula to find the roots of the given function, and then graph the function.
f(x) = x^2 - 4

Answer by stanbon(75887)

f(x) = x^2 - 4
a=1, b=0, c=-4
Using the Quadratic formula you get:
x=-0+sqrt[0^2-4(1)(-4)}/2 = [sqrt(16)]/2 = 4/2=2
OR
x=-0-sqrt[0^2-4(1)(-4)}/2 = -[sqrt(16)]/2 = -4/2= -2

Cheers,
Stan H.